Suppose a perfectly competitive firm with total cost function given as:
TC= 400+20Q-2Q2+2
3Q
3.
A. Find profit maximizing level of output and the maximum profit if the average revenue
equals $180.
B. Calculate the shutdown level of output and price.
1
Expert's answer
2021-08-09T06:47:50-0400

Solution:

A.). A perfectly competitive firm will find its profit-maximizing level of output where MR = MC.

For a perfectly competitive firm: AR = MR = P

AR = 180

Therefore, MR = 180

Derive MC from Total Cost (TC):

TC = 400+20Q-2Q2+23Q3

MC = "\frac{\partial TC} {\partial Q} =" 20 – 4Q + 69Q2 = 69Q2 – 4Q + 20

MC = 69Q2 – 4Q + 20


MC = MR

69Q2 – 4Q + 20 = 180

69Q2 – 4Q + 20 – 180 = 180 – 180

69Q2 – 4Q – 160 = 0

Solve through quadratic function:

Q = 1.55

The profit maximizing level of output = 1.55

 

Profit = TR – TC

TR = P "\times" Q = 180 "\times"1.55 = 279

TC = 400+20Q-2Q2+23Q3

Substitute Quantity in the TC function to derive TC.

TC = 400 + 20(1.55) - 2(1.552) + 23(1.553)

TC = 400 + 31 – 4.81 + 85.65 = 400 + 31 + 85.65 – 4.81 = 511.84

Profit/Loss = TR – TC

Profit/Loss = 279 – 511.84 = (232.84)

Profit/Loss = (232.84)


B.). The shutdown price occurs at the minimum of the average variable (AVC) curve, a point where MC = AVC.

First, derive MC:

MC = "\frac{\partial TC} {\partial Q} =" 20 – 4Q + 69Q2 = 69Q2 – 4Q + 20

MC = 69Q2 – 4Q + 20

Then, derive the AVC:

AVC = "\frac{VC} {Q} =\frac{20Q - 2Q^{2} + 23Q^{3} } { Q} = 20 - 2Q + 23Q^{2}"


Set MC = AVC

69Q2 – 4Q + 20 = 20 – 2Q + 23Q2

69Q2 – 4Q + 20 – 23Q2 + 2Q = 20

46Q2 – 2Q = 0

Solve through quadratic equation:

Q = 0.04

The shutdown level of output = 0.04

Derive shutdown price:

P = MC

P = 69Q2 – 4Q + 20

P = 69(0.04)2 – 4(0.04) + 20

P = 11.04 – 0.16 + 20

P = 19.95

The shutdown price = 19.95